OpenStudy (anonymous):

Evaluate the polynomial for x=7. 3x^2-2x+7

hero (hero):

3(7)^2 - 2(7) + 7 3(49) - 14 + 7 147 - 14 + 7 = 147 - 7 = 140

OpenStudy (anonymous):

140